What is the lowest freezing temperature for a saltwater solution? The solubility of sodium chloride (NaCl) is 280g per 1000g of water at 0°C. kf of H2O = 1.86°C/m.







1
Expert's answer
2021-11-12T15:33:01-0500

M(NaCl) = 58.44 g/mol

n(NaCl) "= \frac{280}{58.44} = 4.79 \;mol"

Molality "m = \frac{4.79}{1 \; kg} = 4.79 \;m"

Dissociation of NaCl is

NaCl → Na+ + Cl-

i = 2 (one mole of Na+ and one mole of Cl-)

"ΔT_f = i \times k_f \times m \\

= 2 \times 1.86 \times 4.79 \\

= 17.82 \; °C \\

ΔT_f = T_f^0 -T_f(sol) \\

T_f(sol) = T_f^0 -ΔT_f \\

= 0 -17.82 \\

= -17.82 \; °C"

Answer: -17.82 °C